OpenStudy (anonymous):

allenbrights electronics is having a 25% off sale. if a radio sale priced at $1.35 what is the regular price

OpenStudy (anonymous):

x*0.75=1.35 x=$1.8

OpenStudy (anonymous):

allenbrights electronics is having a 25% off sale. if a radio sale priced at $135 what is the regular price

OpenStudy (anonymous):

allenbrights electronics is having a 25% off sale. if a radio sale priced at $135 what is the regular price

OpenStudy (anonymous):

let x be the regular price then x-0.25x=$1.35 x(1-0.25)= 1.35 can you solve it for x now???